OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 4 when x = 7. What is the value of x when y = 12?

OpenStudy (anonymous):

Plug it in again: y = kx 4 = k(7) 4 = 7k Divide 7 to both sides, we get 4/7. Now plug that back into the equation: y = kx y = 4/7k Plug in 12: 12 = 4/7k Divide 4/7 to both sides or multiply by 7/4.

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

We're dividing: (12) / (4/7) or multiplying: 12 * 7/4

OpenStudy (anonymous):

is it 21?\

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Thanks!